OpenStudy (anonymous):

If the sum of 3 consecutive natural numbers is equal to 264. Choose the equation that could be used to find the first number, x.

OpenStudy (anonymous):

Let the numbers be X, X+1, X+2 (You get that, right?) So simply X+X+1+X+2 = 264 3X = 264-3 X = 261/3 = 87 :)